OpenStudy (anonymous):

can you help me -I am a mother trying to finish her degree and have an alegbra question

OpenStudy (anonymous):

yes sure

OpenStudy (anonymous):

I need to find a horizontal asymptote in y = 27^2 + 1/x^2 - 1

OpenStudy (anonymous):

y=

OpenStudy (anonymous):

An asymptote, (vertical, horizontal or slant) is a line which is tangent to a curve at infinity.

OpenStudy (anonymous):

Horizontal asymptotes occur when the denominator has an equal or larger leading exponent than the polynomial in the numerator. They indicate general behavior far off to the sides of the graph, and can be found by evaluating tendency of y as x tends to infinity by dividing by highest power in denominator.

OpenStudy (anonymous):

I

OpenStudy (anonymous):

III need gheed help with how to find it!

OpenStudy (anonymous):

we have to differentiate the expression with respect to x and find the solution for dy/dx=0

OpenStudy (anonymous):

so when we do it we get dy/dx=-2x^3 =0 hence x=0 is the answer.

OpenStudy (anonymous):

sorry wait for 2 mins please

OpenStudy (anonymous):

when we differentiate we get -2x/(x^2-1)=0 on solving for x we get x=0 hence answer is x=0

OpenStudy (anonymous):

got it?

OpenStudy (anonymous):

\[y = \frac{27x^2 + 1}{x^2 - 1}\]Is this the problem?

OpenStudy (anonymous):

I will proceed assuming it is. The horizontal asymptote in this case will be the quotient as x gets large in either direction. Note that both the numerator and the denomimator get really large some distance from the origin, so the lower degree terms (in this case plus and minus one) don't really change the size of either the numerator or denominator a long way from the origin. This means the numerator will be about 27 times the size of the denominator at large values of x. This, in turn, indicates a horizontal asymptote at y=27.